Circular No. 5633

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1992 QB_{1}

Further observations, made at the Anglo-Australian Observatory on Sept. 21 and at the European Southern Observatory on Sept. 27, are given in the October batch of Minor Planet Circulars, together with minor corrections to the times of the earlier Mauna Kea observations. The ephemeris below, by B. G. Marsden, Center for Astrophysics, is from a new circular orbit, which differs only marginally from the one on IAUC 5611. MPC 20937 gives a very slightly different ephemeris, putting the object currently about 3.5 AU farther away. That ephemeris is from some hypothetical elliptical elements, given on MPC 20935, computed on the assumption that the object is now near aphelion and similar to (2060) Chiron and (5145) Pholus, with q = 8.5 AU and P ∼ 140 yr.

1992TT α_{2000} δ_{2000} Δ r ϵ β V
Oct. 5 23^{h}58.^{m}58 − 0^{o}07.^{′}9 40.192 41.169 167.^{o}5 0.^{o}3 23.5
15 23 57.86 − 0 12.5 40.246 41.169 157.4 0.5 23.5
25 23 57.21 − 0 16.6 40.328 41.169 147.3 0.7 23.5
Nov. 4 23 56.64 − 0 20.2 40.436 41.169 137.2 0.9 23.6
14 23 56.18 − 0 23.1 40.566 41.169 127.0 1.1 23.6
24 23 55.85 − 0 25.1 40.714 41.169 116.8 1.2 23.6
Dec. 4 23 55.67 − 0 26.3 40.876 41.169 106.6 1.3 23.6
14 23 55.63 − 0 26.4 41.046 41.169  96.5 1.4 23.6
24 23 55.74 − 0 25.6 41.220 41.169  86.3 1.4 23.6

SUPERNOVA 1992aw IN ANONYMOUS GALAXY D. D. Balam, University of Victoria, reports the following precise po-

sition of this object (cf. IAUC 5606), from images obtained by F. D. A. Hartwick, G. C. L. Aikman, and Balam at the 1.85-m telescope of the Dominion Astrophysical Observatory: α = 19^{h}05^{m}40^{s}.48, δ = +50^{o}58′09′′.2 (equinox 1950.0). The center of the host galaxy has end figures 39^{s}.11, 04′′.2.

NOVA SAGITTARII 1992 No. 2 Additional photometry (cf. IAUC 5576) by A. C. Gilmore, Mt. John

University Observatory (± 0.02 unless otherwise noted): Sept. 17.44 UT, V = 11.86, B–V = +0.07, U–B = –0.56, V–R = +1.49 ± 0.04, V–I = +0.68; 29.47, 12.17, +0.11, –0.39, +1.31 ± 0.03, +0.53 ± 0.05.

1992 October 8 Daniel W. E. Green